Mr. Taylor gave his class a test the day after spring break. The average score for all students was 80. The average score for those who passed was 84 and for those that failed was 60. If there were 60 students in the class, how many failed?
Thank you for your help

Answer by stanbon(75887)

Mr. Taylor gave his class a test the day after spring break. The average score for all students was 80. The average score for those who passed was 84 and for those that failed was 60. If there were 60 students in the class, how many failed?
------------------
Equations:
P + F = 60
84P + 60F = 80*60
-------------------
Modify to solve for F::
84P + 84F = 84*60
84P + 60F = 80*60
-----------------------
24F = 4*60
F = 10 (# who failed)
-------------
Cheers,
Stan H.
---------